OpenStudy (anonymous):

Circuits - @whpalmer4 @yakeyglee - If I know power for resistor two and the circuit voltage, how do I solve for resistance? Let me get a drawing and what I have so far...

OpenStudy (anonymous):



OpenStudy (anonymous):

Power is related to voltage drop across a resistor and the resistance of that resistor by \(P = V^2/R\) so \(R = V^2/P\).

OpenStudy (anonymous):

P2=V2*I therefore P2=(R2/(8)R2)I

OpenStudy (whpalmer4):

You can also use \[P=VI\]\[V = IR\]\[P=I*IR = I^2R\] to find the resistance from the power because you know \[I=\frac{V}{8R_2}\]

OpenStudy (anonymous):

I'm sorry @whpalmer4... I still can't get it. \[P=V_{2}I\] but when we get down to the I=V/8R2 i cant find I because i don't know what r2 is?

OpenStudy (anonymous):

I'm so confused

OpenStudy (whpalmer4):

Okay, we know that \[V=IR\]so\[I=\frac{50}{4R_2+R_2+3R_2} = \frac{25}{4R_2}\]We also know that \[P = I^2R_2 = 160 \text{mW}\]Combine the two equations and you should be able to solve for \(R_2\)

OpenStudy (anonymous):

\[.16=\left(\begin{matrix}50 \over\ 8R_{2}\end{matrix}\right)^2 \times R_{2}\]\] \[.16=\left(\begin{matrix}2500 \over\ (64R_{2})^2\end{matrix}\right) \times R_{2}\] \[.16=\left(\begin{matrix}2500 \over\ 64R_{2}\end{matrix}\right)\]

OpenStudy (anonymous):

= .004096 ?

OpenStudy (whpalmer4):

\[(\frac{25}{4R_2})^2R_2=160\text{mW}\]\[\frac{625}{16R_2}=160\text{mW}\]\[R_2=\frac{625}{16*160*10^{-3}} \approx 244.141 \Omega\]\[I=\frac{25}{4*244.141}=25.6\text{mA}\]\[V=I*(R_1+R_2+R_3) =25.6*10^{-3}*(8*244.141) = 50\]

OpenStudy (anonymous):

Wowzers i messed that up quite a bit Let me take in your math so i can actually use this effectively

OpenStudy (whpalmer4):

You mistyped the denominator in the second equation, but didn't actually heed what it said, so got the right 3rd equation. \[0.16=\frac{2500}{64R_2}\]\[0.16*64*R_2=2500\]\[R_2=\frac{2500}{0.16*64}=\frac{2500}{10.24} = 244.140625\]Maybe you just pushed the buttons wrong on the calculator? Or your slide rule needs lubricating :-)

OpenStudy (anonymous):

That is what i was just about to say, when i plugged it into the calc it was a miss type from the start... thank you for clearing this up. looks like this was the hardest problem from the lot, i was able to finish most of them in the time it took working on these two problems with you. Thank you very much @whpalmer4 !!!

OpenStudy (whpalmer4):

You're welcome, of course! What class is this for?

OpenStudy (anonymous):

Ha, embarrassing enough, this is circuits I, which is just a simple 1140 series class...

OpenStudy (whpalmer4):

What's the textbook you are using?

OpenStudy (whpalmer4):

Are you in agreement that the text provides "a thorough, engaging introduction to the theory"? :-)

OpenStudy (whpalmer4):

I think they write that in the review of every circuit theory textbook ever written. I'm not sure it has ever been true :-)

OpenStudy (anonymous):

Wellllllll it seemed like that until i started working on this homework, plus it didnt help that the prof. didnt prepare us for this kind of work... even though it is algebra.

OpenStudy (anonymous):

Haha, no it has not, unfortunately.

OpenStudy (whpalmer4):

Well, it could be worse, in physics textbooks they always want you to compute the resistance of grids of interconnected resistors stretching off to infinity and beyond :-)

OpenStudy (anonymous):

Damn, well that gives me a little peek into my future :( haha. I really appreciate the help though!

OpenStudy (whpalmer4):

But, a lot of this is just a matter of seeing some good examples and working a bunch of problems until you get the hang of it.

OpenStudy (anonymous):

Thats what I am really hoping for

OpenStudy (whpalmer4):

Well, shoot me a message if you get stuck on one and no one seems to be helping. Hopefully the reason for that won't be that the problem is impossibly difficult :-)

OpenStudy (anonymous):

Thank you sooo much, I'd hate to bug you but you really know what you're doing!

OpenStudy (whpalmer4):

Don't sweat it. You won't occupy any more of my time than I'm willing to spend here. As long as whoever is using my time is learning something, I'm happy.

OpenStudy (whpalmer4):

And it turns out that explaining things to someone else is a good way to reinforce your own understanding.

OpenStudy (anonymous):

That is so awesome of you. I try and help out others in math when I can, and when I know how to get it right haha, but lately its been working on this and calcII for hours and hours on end

OpenStudy (anonymous):

It does help quite a bit, that is why I enjoy helping others if I get the chance.

OpenStudy (whpalmer4):

Well, time for me to turn in, it's almost 2:30 AM here on the west coast.

OpenStudy (anonymous):

Sitting here at 3:30 AM almost done though. Thank you so much for the help again!